solve for x and y in the system of equations : sinxcos2y = 1 and cosxsin2y = 0

Grade:11

1 Answers

Y RAJYALAKSHMI
9 years ago
By Dividing the second equation by first equation, we have
cos x sin 2y / sin x cos 2y = 0/1
=> cot x tan 2y = 0
=> cotx = 0  => x = pi/2
or tan2y = 0 => y = 0
 
Solution : x = pi/2  & y = 0